A firm buys and resell goods at 25% markup. It gives 1 weeks's credit from the people it buys from and give 2 weeks credit terms to the people it sell to. Sales per year are 5000. What is your average working capital need for the year? Also what is the income statement when financing cost is 10% and tax rate 40%.
Answer #1

Answer:

Sales per year = 5,000

Sales are done with 25% mark up

Hence cost price = 5,000 / (1 + 25%) = 4,000

It gets one week credit from suppliers.

Hence:

Average accounts payable = 4,000 /52 = 76.9231

It gives 2 weeks credit terms to the people it sells to.

Hence:

Average account receivable = (5,000 / 52) * 2 = 192.3077

Hence, assuming no other current assets /liabilities other than AR and AP

Average working capital need for the year = Average account receivable - Average accounts payable = 192.3077 - 76.9231 = 115.38

Average working capital need for the year = 115.38

Interest / Financing cost = 115.38 * 10% = 11.54

Income Statement is as follows:

Income Statement Sales 5000.00 Cost of Goods sold 4000.00 Gross Income Interest Income before Tax988.46 Tax at 40% Net Income